Jamylla and Janylla are on a hike. Both hikers begin a trip by walking, the first leg A: 2.65 km, 140° from their base camp, on their second leg B: 4.77 km, 252° and to complete their final leg C: 3.18 km, 332°. What is the magnitude and direction of their final position from their base camp? Apply polygon method and component method and compare their results.

1
Expert's answer
2022-03-13T18:49:01-0400

"x=x_a+x_b+x_c=-0.696~km,"

"y=y_a+y_b+y_c=-4.326~km,"

"r=\sqrt{x^2+y^2}=4.382~km,"

"\varphi=180°+\arctan\frac yx=261°."